OpenStudy (anonymous):

Help please. Medals are given. Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (ikram002p):

u can see that f(1)=0 and f(2) =1 so the first two turm is 0,1 so a is not correct now find f(3)=f(3-1)+f(3-2)=f(2)+f(1)=1+0=1 so u have now the first three elements 0,1,1 so c is not correct now find f(4)=f(4-1)+f(4-2)=f(3)+f(2)=1+1=2 so u have now the first 4 elements 0,1,1,2 now find f(5)=f(4)+f(3) the same way as before f(5)=2+1=3 so u have now the first 5 elements 0,1,1,2,3

OpenStudy (anonymous):

So, it is B. I have one more question if you do not mind helping. @ikram002p

OpenStudy (ikram002p):

ok lets see

OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Numerical Answers Expected!

OpenStudy (ikram002p):

f(3)=f(1)+f(2)+f(3-1)=2+4+f(2)=2+4+4=10 f(4)=f(1)+f(2)+f(4-1)=2+4+f(3)=2+4+10=18 f(5)=f(1)+f(2)+f(5-1)=2+4+f(4)=2+4+18=24

OpenStudy (anonymous):

okay so do I put all of that for f(5)=..................... ?

OpenStudy (ikram002p):

now what does "Numerical Answers" mean ?!

OpenStudy (anonymous):

ok I know it now.